OpenStudy (anonymous):

What is the absolute value of the number? 3 + i

myininaya (myininaya):

\[|x+yi|=\sqrt{x^2+y^2}\]

OpenStudy (anonymous):

So it's 4?

myininaya (myininaya):

You have |3+1i|. So Replace x with 3 and y with 1, then evaluate \[\sqrt{3^2+1^2}\]

OpenStudy (anonymous):

when I solve I get 3.16. but my choices are: square root 10 4 8 square root 3

OpenStudy (anonymous):

so it would be square root of 10? Thanks :)

myininaya (myininaya):

Yes nine plus one is ten.